Five years ago, Dylan was half as old as Alyssa was then. Ten years from now, Dylan will be as old as Alyssa was five years ago. How old is Alyssa now?
1
Expert's answer
2012-10-17T09:01:37-0400
Let assume that Alyssa now is X years old and Dylan is Y years old:
X-5=2(Y-5)
Y+10=X-5, Y=X-15
X-5=2(X-20)
X-5=2X-40
X=35